OpenStudy (anonymous):

Determine the value of k so that the lines kx-y=-1/3 and 3y=1-6x will not intersect?

OpenStudy (sidsiddhartha):

\[kx-y=-\frac{ 1 }{ 3 }................(1)\]\[6x+3y=1.....................(2)\] can u find out slopes for those two lines?

OpenStudy (sidsiddhartha):

ok use the standard form \[y=mx+c, ~~where~~m=slope.of.the.line\] so for the first equation \[y=kx+(1/3)\]so for this equation slope \[m_1=k\]and for second equation we can write \[3y=1-6x \\y=(1/3)-2x\\y=-2x+(1/3)\] so for this equation slope will be \[m_2=-2\] now,they will never intersect if their slopes are equal so--\[m_1=m_2\\k=-2\] so for k=-2 they will never intersect

OpenStudy (sidsiddhartha):

got this? @jelolopera

OpenStudy (anonymous):

this k= -2 will not intersect

OpenStudy (sidsiddhartha):

you can check your answer is correct or not by putting k=-2 in the first equation then it will become \[-2x-y=-1/3\\6x+3y=1\]and the second equation is also \[6x+3y=1\] so basilcally the two equations are same means they are parallel so they will never intersect

OpenStudy (sidsiddhartha):

yeah :)